# Geometric Approach to Light Rings in Axially Symmetric Spacetimes

### v2

@@ III Geometric Approach to Determine Light Rings
For any stationary and axially symmetric rotational spacetime, the constructed optical geometry is a Randers-Finsler geometry given by expressions (7) and (8). The stationary time $t$ plays the role of arc-length parameter in optical geometry, and the lengths of photon orbits are measured by the stationary time interval between the emission time $t_{A}$ and the received time $t_{B}$
@@
This is the geodesic curvature condition for light rings in axially symmetric rotational spacetimes. Furthermore, by comparing our geometric approach with the conventional effective potential approach, we establish a correspondence for light rings in rotational spacetimes
@@
Flag Curvatures Condition to the Stability of Light Rings, Conjugate Points, and Cartan-Hadamard Theorem: Having derived the geodesic curvature condition for light rings in equations (17) and (21), it is necessary to further explore the stability of these light rings in the equatorial plane. To carry out a geometric analysis of the light ring stability using geometric properties of optical geometry, we resort to the mathematical concept of conjugate points. The existence of conjugate points proves to be of critical importance in distinguishing the stable and unstable light rings. Mathematically, two points $p$ and $q$ are defined to be conjugate points in a manifold, if a convergence of nearby geodesics starting from point $p$ can eventually converge to another point $q$ [102, 103]. Conversely, if two different geodesic curves starting from the point $p$ can not converge at another point, then there are no points conjugate to $p$ in this manifold. In the optical geometry, stable and unstable circular photon orbits have distinct properties, which have nontrivial connections with the conjugate points. Firstly, if photons are perturbed from an unstable circular photon orbit at a given point $p$, they would inevitably move away from this unstable circular orbit (the perturbed photons may eventually escape to infinity or fall into the event horizon produced by black holes). In such cases, the perturbed photons’ orbit and the original unstable circular photon orbit, which belong to a geodesic congruence staring from the same point $p$, can not converge to each other at another point. Consequently, there are no conjugate points in the unstable circular photon orbit, as illustrated in the left part of figure 2. On the other hand, if photons are perturbed from a stable photon orbit, instead of moving away completely, they may travel along other bound photon orbits in the vicinity of this stable circular photon orbit. In such cases, the original stable circular orbit and the nearby bound photon orbits may have intersect points (the points $q_{1}$ and $q_{2}$ in the right part of figure 2, producing conjugate points in this stable circular photon orbit (for instance, $p$ and $q_{1}$, $q_{1}$ and $q_{2}$). Based on the above analysis, we conclude the following nontrivial relation between the stability of circular photon orbits and the existence of conjugate points:
@@
Figure 2: Illustration of the nontrivial connections between the stability of circular photon orbits and the existence of conjugate points in optical geometry. (a) The left panel illustrates the photon beams perturbed from an unstable circular photon orbit at a given point $p$. The perturbed photons would inevitably move away from this unstable circular orbit (escape to infinity or fall into the event horizon produced by black holes), and it is not possible to find another point $q$ conjugate to $p$ in this unstable circular photon orbit. (b) The right panel illustrates the photon beams perturbed from a stable circular photon orbit. In such cases, the perturbed photons may travel along other bound photon orbits near this stable circular photon orbit. There are conjugate points in this stable circular photon orbit (it is easy to observe that $p$ and $q_{1}$ are conjugate points, meanwhile, $q_{1}$ and $q_{2}$ are also conjugate points).
@@
This Cartan-Hadamard theorem in Randers-Finsler optical geometry gives a correspondence on intrinsic flag curvature, existence of conjugate points, and the stability of light rings. For any stable light ring, the existence of conjugate points in this light ring definitely violates the prerequisite of the Cartan-Hadamard theorem, which suggests that the flag curvature in the vicinity of stable light rings must be positive. Conversely, for the unstable light rings, the absence of conjugate points is consistent with the situations described in the Cartan-Hadamard theorem. This suggests that the flag curvature of optical geometry in the vicinity of unstable light rings is always negative  [^3] There are some subtleties in this conclusion, since our assumption is slightly different from the cases described in the Cartan-Hadamard theorem in mathematical literature. Strictly speaking, the Cartan-Hadamard theorem is a theorem relevant to global geometric and topological properties. The prerequisite of the Cartan-Hadamard theorem is that the flag curvature must be non-positive everywhere, and the conclusion is that no conjugate points exist in the entire Finsler manifold. In the present work, we do not care about whether conjugate points exist in the entire equatorial plane of optical geometry (especially the region very far from the light ring positions). What we are really concerned about is the conjugate points located in light rings (or near light rings). So we slightly weaken the prerequisite condition of the Cartan-Hadamard theorem, assuming the flag curvature in the local region near light rings is negative (from this assumption, we also postulate that light ring position is not geometrically flat). However, the readers need not worry about this point, because an equivalence relation between our geometric approach and the conventional effective potential approach presented in the next section could provide a justification for our assumption.. According to the Cartan-Hadamard theorem in 2-dimensional Randers-Finsler optical geometry, the following flag curvature condition for stable and unstable light rings can be proposed:
@@
From this reduced flag curvature expression, we obtain the following flag curvature condition for stable and unstable light rings in axially symmetric rotational spacetimes
@@
For physical spacetimes, $g_{tt}<0$, $g_{rr}>0$, $g_{\phi\phi}>0$ always hold for circular photon orbits, which suggests $\alpha_{rr}=-\frac{g_{rr}}{g_{tt}}>0$ and $\alpha_{\phi\phi}=\frac{g_{t\phi}^{2}-g_{tt}g_{\phi\phi}}{g_{tt}^{2}}>\beta_{\phi}^{2}=\frac{g_{t\phi}^{2}}{g_{tt}^{2}}\geq 0$, so the first two factors outside the square brackets in the expression (25) is always positive. Furthermore, it is natural to think that our flag curvature conditions for stable and unstable light rings in equations (23a-23b) or (26a-26b) may be equivalent to the effective potential conditions (the local minimum and maximum of the effective potential correspond to stable and unstable light rings), since the different approaches determine the same light rings in a gravitational field. In the next section, we give a demonstration of this equivalence relation
@@ IV Equivalence Between Our Geometric Approach and the Conventional Effective Potential Approach
In this section, we give a demonstration of the equivalence between our geometric approach and the conventional effective potential approach. The demonstration process relies on the analysis and simplification of the effective potential, through the utility of conserved quantities in axially symmetric spacetimes. Consider the stationary and axially symmetric rotational spacetime with the general metric form
@@
The parameter $\epsilon$ distinguishes the particle types in a gravitational field: $\epsilon=0$ characterizes massless photon orbits (e.g., photon orbits) and $\epsilon=1$ corresponds to massive particle orbits. It is worth noting that in many references, an additional minus sign is often absorbed into the definition of effective potential, $V_{\text{eff}}(r)\to-V_{\text{eff}}(r)$. However, our choice of effective potential in expression (33) is made such that the local minimum of effective potential $\frac{d^{2}V_{\text{eff}}}{dr^{2}}>0$ always corresponds to stable particle orbits, and the local maximum of effective potential $\frac{d^{2}V_{\text{eff}}}{dr^{2}}<0$ always corresponds to unstable particle orbits.
@@
Therefore, the unit tangent vector norm condition for circular curves in optical geometry can result in the vanishing of the effective potential for photon beams in the spacetime geometry. Eventually, we have proven that our geometric approach and the conventional effective potential approach yield the same well-known equation for angular velocity, which implies the equivalence of these two approaches when analyzing light rings for axially symmetric rotational spacetimes.
@@
Then we present an analysis of the light rings stability, proving the equivalence between our flag curvature criterion in optical geometry and the local maximum (local minimum) criterion of effective potential for stable (unstable) light rings. To achieve this point, it is necessary to compare the analytical expressions for the second-order derivative of effective potential and the flag curvature in Randers-Finsler optical geometry. The second-order derivative of the effective potential at the light ring positions can be calculated as
@@
Furthermore, the local maximum (or local minimum) criterion for stable (or unstable) photon spheres in spherically symmetric spacetimes is equivalent to the Gaussian curvature criterion for stable (or unstable) photon spheres
@@
Through the relations presented in expressions (38-39), (40), (47), (52a-52b), and (59a-59b), we successfully demonstrate the equivalence between our geometric approach (based on intrinsic curvatures in Randers-Finsler optical geometry) and the conventional approach (based on the effective potential of photons). Given this equivalence, it is reasonable to apply this geometric approach to an arbitrary stationary and axially symmetric rotational spacetime. Our geometric approach, which provides a mathematically self-contained framework for the studies of photon spheres and light rings, serves as an alternative and compliment approach to the conventional effective potential approach. A systematic comparison of our geometric approach to photon spheres in spherically symmetric spacetimes, our geometric approach to light rings in axially symmetric spacetimes, and the conventional effective potential approach is summarized in table 1.
@@ V Some Examples
In this section, we present two representative examples to illustrate how to obtain light rings in axially symmetric spacetimes using our geometric approach. In gravity theories, the most fundamental axially symmetric spacetimes are the Kerr spacetime and Kerr-Newman spacetime. They have profound significance in both theoretical and observational studies. The application of our geometric approach to such spacetimes vividly illustrates the usability and validity of our approach.
@@
This gives the correct light ring positions in Kerr spacetime, derived solely from the geodesic curvature condition in Randers-Finsler optical geometry. It is in full agreement with the well-known analytical results for light rings in Kerr spacetime [104, 105]. Here, we have conveniently defined the direction of increasing azimuthal angle $\phi$ such that a positive angular velocity $\Omega>0$ describes the prograde motion of light, while a negative angular velocity $\Omega<0$ describes the retrograde motion of light. In this convention, the spin parameter of the Kerr spacetime is assumed to be nonnegative $a\geq 0$, with the minus and plus signs in the last equality of expression (64) corresponding to light rings for prograde and retrograde photon motions. Remarkably, the second last line in expression (64) is precisely the constraint that can be obtained from the vanishing of the Carter constant in the equatorial plane of Kerr spacetime [106, 107]
@@
Kerr-Newman Spacetime: The metric for Kerr-Newman spacetime is expressed as
@@
The analytical expression of light ring radius in the Kerr-Newman spacetime is much more complicated than that in Kerr spacetime, which is not presented here. The detailed analytical results on light rings’ radii in the Kerr-Newman spacetime can be consulted in references [108, 109].
@@
To summarize, the two illustrative examples (Kerr spacetime and Kerr-Newman spacetime) presented in this section demonstrate the efficacy of our geometric approach in such axially symmetric rotational spacetimes. The results on the radii of light rings obtained from our approach are in complete agreement with those calculated from the conventional effective potential approach. In principle, our approach can be universally applied to any stationary and axially symmetric spacetime (for the study of light rings in the equatorial plane), regardless of the specified metric form. Due to space limitations, other examples are not given in this section.
VI Conclusion
@@
Circular photon orbits are of paramount significance in the astrophysical observations of black holes and other ultra-compact objects, and they also bring valuable insights into understanding the topological properties of spacetimes, gravitational waves, and quasi-normal modes.
@@
In this work, we present a novel geometric approach to light rings in axially symmetric rotational spacetimes, studying light rings and their stability through intrinsic curvatures in optical geometry. This work provides an extension of our geometric approach to circular photon orbits, which was first proposed in Phys. Rev. D 106, L021501 (2022), from spherically symmetric spacetimes to axially symmetric spacetimes. When extended to axially symmetric spacetimes, the construction of optical geometry gives rise to a Randers-Finsler geometry $dt=\sqrt{\alpha_{ij}dx^{i}dx^{j}}+\beta_{i}dx^{i}$. The location of light rings can be determined by vanishing the geodesic curvature in the equatorial plane of Randers-Finsler optical geometry. Moreover, the stability of light rings can be distinguished through the existence of conjugate points, which is dominantly controlled by the sign of intrinsic flag curvature in Randers-Finsler optical geometry. Specifically, the positive flag curvature indicates a stable light ring, while the negative flag curvature implies an unstable light ring. To validate the effectiveness of our geometric approach, we have chosen two representative examples: Kerr spacetime and Kerr-Newman spacetime. In both cases, our method completely reproduces the correct light ring positions. Beyond these examples, our approach can be applied to an arbitrary stationary spacetime, without any restriction on the specific spacetime metric forms. Furthermore, this work also offers a rigorous demonstration of the equivalence between our geometric approach and the conventional approach based on the effective potential of photons, showing the broad applicability and robustness of our geometric approach. This equivalence relationship, which has been summarized in table 1, not only provides new perspectives for light rings but also builds bridges between different theoretical methods.
Our present work, together with a series of investigations given by the authors and other scholars in recent years, reveals a nontrivial relationship between circular particle orbits and the intrinsic curvatures of the low-dimensional geometry that is constructed from Lorentz spacetime (such as optical geometry and Jacobi geometry). These explorations uncover the deep underlying connections between kinematic characteristics of particle orbits (such as stability of circular orbits, deflection angles) and the intrinsic geometric properties of the low-dimensional constructed geometry of spacetime, indicating the remarkable potential of such mathematical constructed geometry in the gravity theory and mathematical physics.
@@
There are a number of topics that can be inspired by the present work. For instance, our geometric approach can be extended to study the massive particle’s circular orbits in axially symmetric spacetimes, especially the innermost circular orbits (ISCO). To achieve this point, the mathematical construction of Jacobi geometry (rather than the optical geometry) is required [114, 115, 116, 117, 118, 119]. Furthermore, it would be of great importance to conduct a comprehensive analysis on the properties of light rings through our geometric approach, questioning whether a geometric analysis from optical geometry can rederive some important theorems on circular photon orbits for axially symmetric rotational spacetimes reported in literature [22, 24, 23, 25, 32], especially the number of stable and unstable light rings in various class of rotational spacetimes (which is relevant to the topological invariant in these spacetimes).
@@
The appendices present the mathematical preliminaries required for our geometric analysis of circular photon orbits in axially symmetric rotational spacetimes. Appendix A offers a brief introduction to the geodesic equation in Randers-Finsler geometry. Appendix B provides discussions on the mathematical definition of geodesic curvature for continuous curves in Randers-Finsler optical geometry. The decomposition of geodesic curvature in Randers-Finsler geometry $\kappa_{g}^{(F)}=\kappa_{g}^{(\alpha)}+\kappa_{\beta}^{(\alpha)}$, which consists of the geodesic curvature with Riemannian part $\alpha$ and the additional contribution from the non-Riemannian part $\beta$, is presented in this section. Appendix C introduces one of the most important intrinsic curvatures in Randers-Finsler optical geometry — the flag curvature.

### v3

@@ III Geometric Approach to Determine Light Rings
For any stationary and axially symmetric spacetime, the constructed optical geometry is a Randers-Finsler geometry given by expressions (7) and (8). The stationary time $t$ plays the role of arc-length parameter in optical geometry, and the lengths of photon orbits are measured by the stationary time interval between the emission time $t_{A}$ and the received time $t_{B}$
@@
This is the geodesic curvature condition for light rings in axially symmetric spacetimes. Furthermore, by comparing our geometric approach with the conventional effective potential approach, we establish a correspondence for light rings in such rotational spacetimes
@@
Figure 2: Illustration of the nontrivial connections between the stability of circular photon orbits and the existence of conjugate points in optical geometry. (a) The left panel illustrates the photon beams perturbed from an unstable circular photon orbit at a given point $p$. The perturbed photons would inevitably move away from this unstable circular orbit (escape to infinity or fall into the event horizon produced by black holes), and it is not possible to find another point $q$ conjugate to $p$ in this unstable circular photon orbit. (b) The right panel illustrates the photon beams perturbed from a stable circular photon orbit. In such cases, the perturbed photons may travel along other bound photon orbits near this stable circular photon orbit. There are conjugate points in this stable circular photon orbit (it is easy to observe that $p$ and $q_{1}$ are conjugate points; meanwhile, $q_{1}$ and $q_{2}$ are also conjugate points).
Flag Curvatures Condition to the Stability of Light Rings, Conjugate Points, and Cartan-Hadamard Theorem: Having derived the geodesic curvature condition for light rings in equations (17) and (21), it is necessary to further explore the stability of these light rings in the equatorial plane. To carry out a geometric analysis of the light ring stability using geometric properties of optical geometry, we resort to the mathematical concept of conjugate points. The existence of conjugate points turns out to be crucial for distinguishing the stable and unstable light rings. Mathematically, two points $p$ and $q$ are defined to be conjugate points in a manifold, if a convergence of nearby geodesics starting from point $p$ can eventually converge to another point $q$ [103, 104]. Conversely, if two different geodesic curves starting from the point $p$ can not converge at another point, then there are no points conjugate to $p$ in this manifold. In the optical geometry, the properties of stable and unstable circular photon orbits differ significantly, which have nontrivial connections with the conjugate points. Firstly, if photons are perturbed from an unstable circular photon orbit at a given point $p$, they would inevitably move away from this unstable circular orbit (the perturbed photons may eventually escape to infinity or fall into the event horizon produced by black holes). In such cases, the perturbed photons’ orbit and the original unstable circular photon orbit, which belong to a geodesic congruence starting from the same point $p$, can not converge to each other at another point. Consequently, there are no conjugate points in the unstable circular photon orbit, as illustrated in the left part of Figure 2. On the other hand, if photons are perturbed from a stable photon orbit, instead of moving away completely, they may travel along other bound photon orbits in the vicinity of this stable circular photon orbit. In such cases, the original stable circular orbit and the nearby bound photon orbits may have intersect points (the points $q_{1}$ and $q_{2}$ in the right part of Figure 2, producing conjugate points in this stable circular photon orbit (for instance, $p$ and $q_{1}$, $q_{1}$ and $q_{2}$). Based on the above analysis, we conclude the following nontrivial relation between the stability of circular photon orbits and the existence of conjugate points:
@@
This Cartan-Hadamard theorem in Randers-Finsler optical geometry gives a correspondence among the intrinsic flag curvature, the existence of conjugate points, and the stability of light rings. For any stable light ring, the existence of conjugate points in this light ring definitely violates the prerequisite of the Cartan-Hadamard theorem, which suggests that the flag curvature in the vicinity of stable light rings must be positive. Conversely, for the unstable light rings, the absence of conjugate points is consistent with the cases described in the Cartan-Hadamard theorem. This suggests that the flag curvature in the vicinity of unstable light rings is always negative  [^3] There are some subtleties in this conclusion, since our assumption is slightly different from the cases described in the Cartan-Hadamard theorem in mathematical literature. Strictly speaking, the Cartan-Hadamard theorem is a theorem relevant to global geometric and topological properties. The prerequisite of the Cartan-Hadamard theorem is that the flag curvature must be non-positive everywhere, and the conclusion is that no conjugate points exist in the entire Finsler manifold. In the present work, we do not care about whether conjugate points exist in the entire equatorial plane of optical geometry (especially the region very far from the light ring positions). What we are really concerned about is the conjugate points located in light rings (or near the light rings). So we slightly weaken the prerequisite condition of the Cartan-Hadamard theorem, assuming the flag curvature in the local region near light rings is negative (from this assumption, we also postulate that light ring position is not geometrically flat). However, the readers need not worry about this point, because an equivalence relation between our geometric approach and the conventional effective potential approach presented in the next section could provide a justification for our assumption.. According to the Cartan-Hadamard theorem in 2-dimensional Randers-Finsler optical geometry, the following flag curvature condition for stable and unstable light rings can be proposed:
@@
From this reduced flag curvature expression, we obtain the following flag curvature condition for stable and unstable light rings in axially symmetric spacetimes
@@
For physical spacetimes, $g_{tt}<0$, $g_{rr}>0$, $g_{\phi\phi}>0$ always hold for circular photon orbits, which suggests $\alpha_{rr}=-\frac{g_{rr}}{g_{tt}}>0$ and $\alpha_{\phi\phi}=\frac{g_{t\phi}^{2}-g_{tt}g_{\phi\phi}}{g_{tt}^{2}}>\beta_{\phi}^{2}=\frac{g_{t\phi}^{2}}{g_{tt}^{2}}\geq 0$, so the first two factors outside the square brackets in the expression (25) are always positive. Furthermore, it is natural to think that our flag curvature conditions for stable and unstable light rings in equations (23a-23b) or (26a-26b) may be equivalent to the effective potential conditions (the local minima and local maxima of the effective potential correspond to stable and unstable light rings), since the different approaches determine the same light rings in a gravitational field. In the next section, we give a demonstration of this equivalence relation
@@ IV Equivalence Between Our Geometric Approach and the Conventional Effective Potential Approach
In this section, we give a demonstration of the equivalence between our geometric approach and the conventional effective potential approach. The demonstration process relies on the analysis and simplification of the effective potential, through the utility of conserved quantities in axially symmetric spacetimes. Consider the stationary and axially symmetric spacetime with the general metric form
@@
The parameter $\epsilon$ distinguishes the particle types in a gravitational field: $\epsilon=0$ characterizes massless photon orbits (e.g., photon orbits) and $\epsilon=1$ corresponds to massive particle orbits. It is worth noting that in many references, an additional minus sign is often absorbed into the definition of effective potential, $V_{\text{eff}}(r)\to-V_{\text{eff}}(r)$. However, our choice of effective potential in expression (33) is made such that the local minima of effective potential $\frac{d^{2}V_{\text{eff}}}{dr^{2}}>0$ always correspond to stable particle orbits, and the local maxima of effective potential $\frac{d^{2}V_{\text{eff}}}{dr^{2}}<0$ always correspond to unstable particle orbits.
@@
Therefore, the unit tangent vector norm condition for circular curves in optical geometry can result in the vanishing of the effective potential for photon beams in the spacetime geometry. Eventually, we have proven that our geometric approach and the conventional effective potential approach yield the same well-known equation for angular velocity, which implies the equivalence of these two approaches when analyzing light rings in axially symmetric spacetimes.
@@
Then we present an analysis of the light ring’s stability, proving the equivalence between our flag curvature criterion in optical geometry and the local maxima (local minima) criterion of effective potential for stable (unstable) light rings. To achieve this point, it is necessary to compare the analytical expressions for the second-order derivative of effective potential and the flag curvature in Randers-Finsler optical geometry. The second-order derivative of the effective potential at the light ring positions can be calculated as
@@
Furthermore, the local maxima (or local minima) criterion for stable (or unstable) photon spheres in spherically symmetric spacetimes is equivalent to the Gaussian curvature criterion for stable (or unstable) photon spheres
@@
Through the relations presented in expressions (38-39), (40), (47), (52a-52b), and (59a-59b), we successfully demonstrate the equivalence between our geometric approach (based on intrinsic curvatures in Randers-Finsler optical geometry) and the conventional approach (based on the effective potential of photons). Given this equivalence, it is reasonable to apply this geometric approach to an arbitrary axially symmetric spacetime. Our geometric approach, which provides a mathematically self-contained framework in the studies of photon spheres and light rings, serves as an alternative and complementary approach to the conventional effective potential approach. A systematic comparison of our geometric approach to photon spheres in spherically symmetric spacetimes, our geometric approach to light rings in axially symmetric spacetimes, and the conventional effective potential approach is summarized in Table 1.
@@ V Some Examples
In this section, we present two representative examples to illustrate how to obtain light rings in axially symmetric spacetimes using our geometric approach. In gravity theories, the most fundamental axially symmetric spacetimes are the Kerr spacetime and Kerr-Newman spacetime. They have profound significance in both theoretical and observational studies. The application of our geometric approach to such spacetimes effectively demonstrates its practical utility and theoretical validity.
@@
This gives the correct positions of light rings in Kerr spacetime, derived solely from the geodesic curvature condition in Randers-Finsler optical geometry. It is in full agreement with the well-known analytical results for light rings in Kerr spacetime [105, 106]. Here, we have conveniently defined the direction of increasing azimuthal angle $\phi$ such that a positive angular velocity $\Omega>0$ describes the prograde motion of light, and a negative angular velocity $\Omega<0$ describes the retrograde motion of light. In this convention, the spin parameter of the Kerr spacetime is assumed to be nonnegative $a\geq 0$, with the minus and plus signs in the last equality of expression (64) corresponding to light rings for prograde and retrograde photon motions. Remarkably, the second last line in expression (64) is precisely the constraint that can be obtained from the vanishing of the Carter constant in the equatorial plane of Kerr spacetime [107, 108]
@@
Kerr-Newman Spacetime (with Electric and Magnetic Charge): The metric for Kerr-Newman spacetime is expressed as
@@
The analytical expression of light ring radii in the Kerr-Newman spacetime is much more complicated than that in Kerr spacetime, which is not presented here. The detailed analytical results on light ring radii in the Kerr-Newman spacetime can be consulted in references [109, 110].
@@
To summarize, the two illustrative examples (Kerr spacetime and Kerr-Newman spacetime) presented in this section demonstrate the efficacy of our geometric approach in such axially symmetric spacetimes. The results on the radii of light rings obtained from our approach are in complete agreement with those calculated from the conventional effective potential approach. In principle, our approach can be universally applied to any stationary and axially symmetric spacetime (for the study of light rings in the equatorial plane), regardless of the specified metric form. Due to space limitations, other examples are not given in this section.
VI Conclusion and Perspectives
@@
Circular photon orbits are of paramount significance in astrophysical observations of black holes and other ultra-compact objects, and they also bring valuable insights into understanding the topological properties of spacetimes, gravitational waves, and quasi-normal modes.
In this work, we present a novel geometric approach to light rings in axially symmetric spacetimes, studying light rings and their stability through intrinsic curvatures in optical geometry. This work provides a natural extension of our geometric approach to circular photon orbits (proposed in Phys. Rev. D 106, L021501 (2022)) from spherically symmetric spacetimes to axially symmetric spacetimes.
@@
In axially symmetric spacetimes, the construction of optical geometry gives rise to a Randers-Finsler geometry $dt=\sqrt{\alpha_{ij}dx^{i}dx^{j}}+\beta_{i}dx^{i}$. Our analysis shows that the location of light rings can be determined by vanishing the geodesic curvature in the equatorial plane of Randers-Finsler optical geometry. Moreover, the stability of light rings can be distinguished through the existence of conjugate points, which are dominantly controlled by the sign of intrinsic flag curvature in Randers-Finsler optical geometry. Specifically, the positive flag curvature indicates a stable light ring, while the negative flag curvature implies an unstable light ring. To validate the effectiveness of our geometric approach, we have chosen two representative examples: Kerr spacetime and Kerr-Newman spacetime. In both instances, our method completely reproduces the correct light ring positions. Generally speaking, our approach can be applied to an arbitrary stationary spacetime, without any restrictions on the specific forms of the spacetime metrics. Furthermore, we provide a rigorous demonstration of the equivalence between our geometric approach and the conventional approach based on the effective potential of photons, showing the broad applicability and robustness of our geometric approach. This equivalence relationship, as summarized in Table 1, not only provides novel perspectives on light rings but also builds bridges between different theoretical methods.
@@
The present work can inspire several research topics. For instance, our geometric approach can be extended to study the circular orbits of massive particles in axially symmetric spacetimes, particularly the innermost stable circular orbits (ISCO). To achieve this point, the mathematical construction of Jacobi geometry (rather than the optical geometry) is necessary [115, 116, 117, 118, 119, 120]. Furthermore, it is highly significant to conduct a comprehensive analysis of the general properties of light rings through our geometric approach. This exploration could examine whether a geometric analysis from optical geometry can rederive some well-established theorems on circular photon orbits for axially symmetric spacetimes reported in literature [23, 25, 24, 26, 33], especially those regarding the existence and number of stable and unstable light rings in various classes of rotational spacetimes.
Our present work, combined with a series of investigations by the authors and other researchers in recent years, reveals a nontrivial relationship between circular particle orbits and the intrinsic curvatures of the low-dimensional geometry constructed from Lorentz spacetime (e.g., optical geometry and Jacobi geometry). The explorations in this direction are likely to uncover deeper and intricate connections between particle orbits in gravitational fields and the geometric properties of these low-dimensional constructed geometries, showing the significant potential of such geometrization methods in gravitational theory and mathematical physics.
@@
The appendices present the mathematical preliminaries required for our geometric analysis of circular photon orbits in axially symmetric spacetimes. Appendix A offers a brief introduction to the geodesic equation in Randers-Finsler geometry. Appendix B discusses the mathematical definition of geodesic curvature for continuous curves in Randers-Finsler optical geometry. The decomposition of geodesic curvature in Randers-Finsler geometry $\kappa_{g}^{(F)}=\kappa_{g}^{(\alpha)}+\kappa_{\beta}^{(\alpha)}$, which consists of the geodesic curvature with Riemannian part $\alpha$ and the additional contribution from the non-Riemannian part $\beta$, is presented in this section. Appendix C introduces one of the most important intrinsic curvatures in Randers-Finsler optical geometry — the flag curvature.